OpenStudy (anonymous):

What are the solutions of the inequality? Graph the solutions. x over 5≥−2 a) x≥−7 b) x≥7 c) x≤−10 d) x≥−10

OpenStudy (anonymous):

\[\frac{ x }{ 5 }\ge-2\]

OpenStudy (anonymous):

a)\[x \ge-7\] b)\[x \ge7\] c)\[x \le-10\] d)\[x \ge-10\]

OpenStudy (anonymous):

do you know how to solve for x in an equation?

OpenStudy (anonymous):

I believe it is B but lets work through this

OpenStudy (anonymous):

i have an idea but just need a little help on making sure i got the right one

OpenStudy (anonymous):

plug in each number for x and see if that number is correct

OpenStudy (anonymous):

Plug the numbers in before you answer the question to check it because I might be wrong

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Oh gosh I was really off I think it's D

OpenStudy (anonymous):

i thought i was D but wasnt sure so once u said B i got al worked up tryn to figure out how u got b lol

OpenStudy (anonymous):

Sorry haha that was a first glance assumption, it's definitely D

OpenStudy (anonymous):

THx :D